OpenStudy (anonymous):

simplify: x^2-3x-10/x+2

OpenStudy (anonymous):

\[\frac{x^2-3x-10}{x+2}\] Try factoring the numerator.

Directrix (directrix):

On problems like this, the denominator is frequently a factor of the numerator. So try this: x^2-3x-10 = ( x + 2)* ( ? + ? )

OpenStudy (anonymous):

(x+5)?

OpenStudy (anonymous):

Close - when there are 2 minus signs in the expanded version (x^2-3x-10), then it usually means that the factorisation contains one plus and one minus.

OpenStudy (anonymous):

The foolproof way of checking your answer would be to expand your roots see if it matches. Does (x+2)(x+5)=x^2-3x-10?

OpenStudy (anonymous):

@KendrickLamar2014, yes, you can...

Directrix (directrix):

Correct or not?